\begin{thebibliography}{17}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Adelani et~al.(2021)Adelani, Ruiter, Alabi, Adebonojo, Ayeni, Adeyemi, Awokoya, and Espa{\~n}a-Bonet]{adelani-etal-2021-effect}
David Adelani, Dana Ruiter, Jesujoba Alabi, Damilola Adebonojo, Adesina Ayeni, Mofe Adeyemi, Ayodele~Esther Awokoya, and Cristina Espa{\~n}a-Bonet.
\newblock The effect of domain and diacritics in {Y}oruba{--}{E}nglish neural machine translation.
\newblock In Kevin Duh and Francisco Guzm{\'a}n (eds.), \emph{Proceedings of Machine Translation Summit XVIII: Research Track}, pp.\  61--75, Virtual, August 2021. Association for Machine Translation in the Americas.
\newblock URL \url{https://aclanthology.org/2021.mtsummit-research.6}.

\bibitem[Adelani et~al.(2022)Adelani, Alabi, Fan, Kreutzer, Shen, Reid, Ruiter, Klakow, Nabende, Chang, Gwadabe, Sackey, Dossou, Emezue, Leong, Beukman, Muhammad, Jarso, Yousuf, Niyongabo~Rubungo, Hacheme, Wairagala, Nasir, Ajibade, Ajayi, Gitau, Abbott, Ahmed, Ochieng, Aremu, Ogayo, Mukiibi, Ouoba~Kabore, Kalipe, Mbaye, Tapo, Memdjokam~Koagne, Munkoh-Buabeng, Wagner, Abdulmumin, Awokoya, Buzaaba, Sibanda, Bukula, and Manthalu]{adelani-etal-2022-thousand}
David Adelani, Jesujoba Alabi, Angela Fan, Julia Kreutzer, Xiaoyu Shen, Machel Reid, Dana Ruiter, Dietrich Klakow, Peter Nabende, Ernie Chang, Tajuddeen Gwadabe, Freshia Sackey, Bonaventure F.~P. Dossou, Chris Emezue, Colin Leong, Michael Beukman, Shamsuddeen Muhammad, Guyo Jarso, Oreen Yousuf, Andre Niyongabo~Rubungo, Gilles Hacheme, Eric~Peter Wairagala, Muhammad~Umair Nasir, Benjamin Ajibade, Tunde Ajayi, Yvonne Gitau, Jade Abbott, Mohamed Ahmed, Millicent Ochieng, Anuoluwapo Aremu, Perez Ogayo, Jonathan Mukiibi, Fatoumata Ouoba~Kabore, Godson Kalipe, Derguene Mbaye, Allahsera~Auguste Tapo, Victoire Memdjokam~Koagne, Edwin Munkoh-Buabeng, Valencia Wagner, Idris Abdulmumin, Ayodele Awokoya, Happy Buzaaba, Blessing Sibanda, Andiswa Bukula, and Sam Manthalu.
\newblock A few thousand translations go a long way! leveraging pre-trained models for {A}frican news translation.
\newblock In \emph{Proceedings of the 2022 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies}, pp.\  3053--3070, Seattle, United States, July 2022. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/2022.naacl-main.223}.
\newblock URL \url{https://aclanthology.org/2022.naacl-main.223}.

\bibitem[Agi{\'c} \& Vuli{\'c}(2019)Agi{\'c} and Vuli{\'c}]{agic-vulic-2019-jw300}
{\v{Z}}eljko Agi{\'c} and Ivan Vuli{\'c}.
\newblock {JW}300: A wide-coverage parallel corpus for low-resource languages.
\newblock In Anna Korhonen, David Traum, and Llu{\'\i}s M{\`a}rquez (eds.), \emph{Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics}, pp.\  3204--3210, Florence, Italy, July 2019. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/P19-1310}.
\newblock URL \url{https://aclanthology.org/P19-1310}.

\bibitem[Chung et~al.(2023)Chung, Garcia, Roberts, Tay, Firat, Narang, and Constant]{chung2023unimax}
Hyung~Won Chung, Xavier Garcia, Adam Roberts, Yi~Tay, Orhan Firat, Sharan Narang, and Noah Constant.
\newblock Unimax: Fairer and more effective language sampling for large-scale multilingual pretraining.
\newblock In \emph{The Eleventh International Conference on Learning Representations}, 2023.
\newblock URL \url{https://openreview.net/forum?id=kXwdL1cWOAi}.

\bibitem[Mubarak et~al.(2019)Mubarak, Abdelali, Sajjad, Samih, and Darwish]{mubarak-etal-2019-highly}
Hamdy Mubarak, Ahmed Abdelali, Hassan Sajjad, Younes Samih, and Kareem Darwish.
\newblock Highly effective {A}rabic diacritization using sequence to sequence modeling.
\newblock In Jill Burstein, Christy Doran, and Thamar Solorio (eds.), \emph{Proceedings of the 2019 Conference of the North {A}merican Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers)}, pp.\  2390--2395, Minneapolis, Minnesota, June 2019. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/N19-1248}.
\newblock URL \url{https://aclanthology.org/N19-1248}.

\bibitem[N{\'a}plava et~al.(2018)N{\'a}plava, Straka, Stra{\v{n}}{\'a}k, and Haji{\v{c}}]{naplava-etal-2018-diacritics}
Jakub N{\'a}plava, Milan Straka, Pavel Stra{\v{n}}{\'a}k, and Jan Haji{\v{c}}.
\newblock Diacritics restoration using neural networks.
\newblock In Nicoletta Calzolari, Khalid Choukri, Christopher Cieri, Thierry Declerck, Sara Goggi, Koiti Hasida, Hitoshi Isahara, Bente Maegaard, Joseph Mariani, H{\'e}l{\`e}ne Mazo, Asuncion Moreno, Jan Odijk, Stelios Piperidis, and Takenobu Tokunaga (eds.), \emph{Proceedings of the Eleventh International Conference on Language Resources and Evaluation ({LREC} 2018)}, Miyazaki, Japan, May 2018. European Language Resources Association (ELRA).
\newblock URL \url{https://aclanthology.org/L18-1247}.

\bibitem[N{\'{a}}plava et~al.(2021)N{\'{a}}plava, Straka, and Strakov{\'{a}}]{naplava-etal-2021-diacritics}
Jakub N{\'{a}}plava, Milan Straka, and Jana Strakov{\'{a}}.
\newblock Diacritics restoration using {BERT} with analysis on czech language.
\newblock \emph{CoRR}, abs/2105.11408, 2021.
\newblock URL \url{https://arxiv.org/abs/2105.11408}.

\bibitem[Oladipo et~al.(2023)Oladipo, Adeyemi, Ahia, Owodunni, Ogundepo, Adelani, and Lin]{oladipo-etal-2023-better}
Akintunde Oladipo, Mofetoluwa Adeyemi, Orevaoghene Ahia, Abraham Owodunni, Odunayo Ogundepo, David Adelani, and Jimmy Lin.
\newblock Better quality pre-training data and t5 models for {A}frican languages.
\newblock In Houda Bouamor, Juan Pino, and Kalika Bali (eds.), \emph{Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing}, pp.\  158--168, Singapore, December 2023. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/2023.emnlp-main.11}.
\newblock URL \url{https://aclanthology.org/2023.emnlp-main.11}.

\bibitem[Orife(2018)]{orife2018attentive}
Iroro Orife.
\newblock Attentive sequence-to-sequence learning for diacritic restoration of yor{\`u}b{\'a} language text.
\newblock \emph{Proc. Interspeech 2018}, pp.\  2848--2852, 2018.

\bibitem[Orife et~al.(2020)Orife, Adelani, Fasubaa, Williamson, Oyewusi, Wahab, and Tubosun]{orife2020improving}
Iroro Orife, David~I. Adelani, Timi Fasubaa, Victor Williamson, Wuraola~Fisayo Oyewusi, Olamilekan Wahab, and Kola Tubosun.
\newblock Improving yor\`ub\'a diacritic restoration, 2020.

\bibitem[Popovi{\'c}(2015)]{popovic-2015-chrf}
Maja Popovi{\'c}.
\newblock chr{F}: character n-gram {F}-score for automatic {MT} evaluation.
\newblock In Ond{\v{r}}ej Bojar, Rajan Chatterjee, Christian Federmann, Barry Haddow, Chris Hokamp, Matthias Huck, Varvara Logacheva, and Pavel Pecina (eds.), \emph{Proceedings of the Tenth Workshop on Statistical Machine Translation}, pp.\  392--395, Lisbon, Portugal, September 2015. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/W15-3049}.
\newblock URL \url{https://aclanthology.org/W15-3049}.

\bibitem[Post(2018)]{post-2018-call}
Matt Post.
\newblock A call for clarity in reporting {BLEU} scores.
\newblock In \emph{Proceedings of the Third Conference on Machine Translation: Research Papers}, pp.\  186--191, Belgium, Brussels, October 2018. Association for Computational Linguistics.
\newblock URL \url{https://www.aclweb.org/anthology/W18-6319}.

\bibitem[Raffel et~al.(2020)Raffel, Shazeer, Roberts, Lee, Narang, Matena, Zhou, Li, and Liu]{2020t5}
Colin Raffel, Noam Shazeer, Adam Roberts, Katherine Lee, Sharan Narang, Michael Matena, Yanqi Zhou, Wei Li, and Peter~J. Liu.
\newblock Exploring the limits of transfer learning with a unified text-to-text transformer.
\newblock \emph{Journal of Machine Learning Research}, 21\penalty0 (140):\penalty0 1--67, 2020.
\newblock URL \url{http://jmlr.org/papers/v21/20-074.html}.

\bibitem[Skiredj \& Berrada(2024)Skiredj and Berrada]{skiredj2024arabic}
Abderrahman Skiredj and Ismail Berrada.
\newblock Arabic text diacritization in the age of transfer learning: Token classification is all you need, 2024.

\bibitem[Stankevicius et~al.(2022)Stankevicius, Lukosevicius, Kapociute{-}Dzikiene, Briediene, and Krilavicius]{stankevicius-etal-2022-correcting}
Lukas Stankevicius, Mantas Lukosevicius, Jurgita Kapociute{-}Dzikiene, Monika Briediene, and Tomas Krilavicius.
\newblock Correcting diacritics and typos with byt5 transformer model.
\newblock \emph{CoRR}, abs/2201.13242, 2022.
\newblock URL \url{https://arxiv.org/abs/2201.13242}.

\bibitem[Wolf et~al.(2020)Wolf, Debut, Sanh, Chaumond, Delangue, Moi, Cistac, Rault, Louf, Funtowicz, Davison, Shleifer, von Platen, Ma, Jernite, Plu, Xu, Le~Scao, Gugger, Drame, Lhoest, and Rush]{wolf-etal-2020-transformers}
Thomas Wolf, Lysandre Debut, Victor Sanh, Julien Chaumond, Clement Delangue, Anthony Moi, Pierric Cistac, Tim Rault, Remi Louf, Morgan Funtowicz, Joe Davison, Sam Shleifer, Patrick von Platen, Clara Ma, Yacine Jernite, Julien Plu, Canwen Xu, Teven Le~Scao, Sylvain Gugger, Mariama Drame, Quentin Lhoest, and Alexander Rush.
\newblock Transformers: State-of-the-art natural language processing.
\newblock In Qun Liu and David Schlangen (eds.), \emph{Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing: System Demonstrations}, pp.\  38--45, Online, October 2020. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/2020.emnlp-demos.6}.
\newblock URL \url{https://aclanthology.org/2020.emnlp-demos.6}.

\bibitem[Xue et~al.(2021)Xue, Constant, Roberts, Kale, Al-Rfou, Siddhant, Barua, and Raffel]{xue-etal-2021-mt5}
Linting Xue, Noah Constant, Adam Roberts, Mihir Kale, Rami Al-Rfou, Aditya Siddhant, Aditya Barua, and Colin Raffel.
\newblock m{T}5: A massively multilingual pre-trained text-to-text transformer.
\newblock In Kristina Toutanova, Anna Rumshisky, Luke Zettlemoyer, Dilek Hakkani-Tur, Iz~Beltagy, Steven Bethard, Ryan Cotterell, Tanmoy Chakraborty, and Yichao Zhou (eds.), \emph{Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies}, pp.\  483--498, Online, June 2021. Association for Computational Linguistics.
\newblock \doi{10.18653/v1/2021.naacl-main.41}.
\newblock URL \url{https://aclanthology.org/2021.naacl-main.41}.

\end{thebibliography}

\documentclass{article} 
\usepackage{iclr2024_conference,times}

\usepackage{amsmath,amsfonts,bm}

\def\eqref#1{equation~\ref{#1}}

\def\1{\bm{1}}

\DeclareMathAlphabet{\mathsfit}{\encodingdefault}{\sfdefault}{m}{sl}
\SetMathAlphabet{\mathsfit}{bold}{\encodingdefault}{\sfdefault}{bx}{n}

\usepackage{hyperref}
\usepackage{url}
\usepackage{graphicx}
\usepackage{xspace,mfirstuc,tabulary}
\usepackage{times}
\usepackage{latexsym}
\usepackage{booktabs}
\usepackage{multirow}
\usepackage{comment}
\usepackage{dirtytalk}
\usepackage{color, colortbl}
\usepackage{adjustbox}
\usepackage{todonotes}
\usepackage{lingmacros}
\usepackage{lipsum} 
\usepackage{subcaption}
\usepackage{cleveref}

\newcommand*{\yoruba}{Yor\`ub\'a\xspace}
\newcommand*{\menyo}{MENYO-20k\xspace}

\definecolor{GrayColor}{gray}{0.9}

\title{YAD: Leveraging T5 for improved automatic diacritization of \yoruba text}

\author{Akindele Michael Olawole\textsuperscript{1}\thanks{Equal contribution.}, Jesujoba O. Alabi\textsuperscript{2*}, Aderonke Busayo Sakpere\textsuperscript{1}, David I. Adelani\textsuperscript{3} \\
\textsuperscript{1} Department of Computer Science, University of Ibadan, Nigeria  \\
\textsuperscript{2} Spoken Language Systems (LSV), Saarland University,
Saarland Informatics Campus, Germany\\
\textsuperscript{3} University College London \\
\texttt{jalabi@lsv.uni-saarland.de}, \texttt{d.adelani@ucl.ac.uk} \\
}

\iclrfinalcopy 
\begin{document}

\maketitle

\begin{abstract}
In this work, we present Yorùbá automatic diacritization (YAD) benchmark dataset for evaluating \yoruba diacritization systems. In addition, we pre-train text-to-text transformer, T5 model for \yoruba and showed that this model outperform several multilingually trained T5 models. Lastly, we showed that more data and larger models are better at diacritization for \yoruba
\end{abstract}

\paragraph{Introduction}
\yoruba, a language spoken predominantly in West Africa, is renowned for its tonal nature which is characterized by a heavy use of diacritics to signify tone variations. In \yoruba and many other languages, diacritics play a crucial role in disambiguating word meanings and in word pronunciation, making accurate diacritization essential for effective communication and language processing tasks~\citep{skiredj2024arabic}. However, manual diacritization is time-consuming and requires specialized linguistic expertise, motivating the development of automatic diacritization systems. In recent years, significant progress has been made in natural language processing (NLP) techniques, leading to the exploration of various approaches to automate the diacritization process for languages using diacritics~\cite[\textit{inter alia}]{naplava-etal-2018-diacritics, mubarak-etal-2019-highly, naplava-etal-2021-diacritics, stankevicius-etal-2022-correcting} including \yoruba~\citep{orife2018attentive, orife2020improving}. Despite these efforts, there is the absence of a standard benchmark dataset for evaluating and comparing developed \yoruba diacritization systems. Furthermore, the emphasis has predominantly been on model development rather than assessing the usability of these systems, particularly in the context of lightweight models amidst the prevalence of large neural networks in contemporary NLP research. In this study, we address these gaps by introducing YAD (\yoruba Automatic Diacritization Dataset), which we curated by leveraging the \menyo~\citep{adelani-etal-2021-effect} dataset. Moreover, we focus on developing lightweight diacritizers for Yoruba, leveraging T5~\citep{2020t5} a text-to-text transformer architecture, to improve usability and efficiency in the era of large neural networks. For reproducibility, we release our code and data on GitHub.\footnote{\url{https://github.com/ajesujoba/YAD}}

\paragraph{\yoruba language} \yoruba orthography is based on the Latin alphabet. Its alphabets include 25 Latin letters excluding characters c, q, v, x and z, and its has five additional letter ({\d e}, gb, {\d s} , {\d o}) representing certain sounds. Diacritics including accents and undots are used in \yoruba to capture tone and vowel height. The tones include low, middle and high which are denoted by the grave (e.g. ``\`{a}''), optional macron (e.g. ``\={a}'') and acute (e.g. ``\'{a}'') accents respectively. 

\paragraph{\yoruba automatic diacritization dataset (YAD)} 
In this work, we present YAD, a dataset created for evaluating diacritization systems for \yoruba. YAD is constructed based on the \menyo dataset, which is a benchmark for English-\yoruba translation, containing 20,000 parallel sentences. We divided the \yoruba side of \menyo's test data into two halves, allocating one half as development set (Dev.) and the other half as test set (Test). We ensured that both halves of the data had equal representation across all domains present in \menyo. Given these splits, we applied four heuristics on the data, considering the sequence-to-sequence nature of the diacritization task. This task requires a source side and a target side, where the source side can consist of \yoruba text without diacritics or partially diacritized texts, and the target side is fully diacritized \yoruba texts. To create and prepare the source side of the data, the following heuristics were applied with specific proportions on the original, well-diacritized texts: (i) removal of all diacritics for $60\%$ of the data, (ii) removal of only tone marks for $20\%$ of the data, (iii) a combination of (ii) and the removal of any random word for $10\%$, and (iv) a combination of (ii) and swapping any two distinct words (applicable only for sentences with more than two words). Where (iii) and (iv) can be seen as infilling and grammatical error correction tasks respectively. Example~\ref{ex:example1}, shows an example from the development set where (iv) was applied where sib{\d e}sib{\d e}, and {\d S}ugb{\d o}n were swapped on the source side.

\enumsentence{\label{ex:example1}\small\textbf{source}: sib{\d e}sib{\d e}, {\d S}ugb{\d o}n Mama o gbagb{\d o}. \\
\textbf{reference}: {\d S}{\` u}gb{\d {\' u}}n s{\' i}b{\d {\` e}}s{\' i}b{\d {\` e}}, M{\` a}m{\' a} {\` o} gb{\` a}gb{\d {\' o}}}

\paragraph{Diacritization model} We pre-train T5 models from scratch using \yoruba WURA corpus~\citep{oladipo-etal-2023-better} and JW300~\citep{agic-vulic-2019-jw300} with different sizes: tiny (number of layers encoder/decoder (L=4), heads (H=4)) , mini (L=4, H=4), small (L=8, H=6), base (L=12, H=12) corresponding to 14M, 18M, 60M and 280M parameters. All models make use of a vocabulary size of 32K except the tiny model. We called our model o\textbf{yo}-T5.~\footnote{Oyo \yoruba (YO) dialect of T5. Oyo dialect is the standardized dialect for writing \yoruba}. All models were pre-trained using HuggingFace Flax code and TPU v3-8 for 100-200K steps, which took around one day each. 

After pre-training, we trained our baseline diacritizer models by fine-tuning existing multilingually trained language models, MT5~\citep{xue-etal-2021-mt5}, AfriMT5~\citep{adelani-etal-2022-thousand}, AfriTeVa-V2~\citep{oladipo-etal-2023-better}, and UMT5~\citep{chung2023unimax}. We use MT5-base,  Afri-MT5-base, the base and large variants of AfriTeVa-V2, and UMT5. we fine-tune these models using HuggingFace Transformers~\citep{wolf-etal-2020-transformers}. These eventual diacritization models are evaluated using the test and development SacreBLEU~\citep{post-2018-call}\footnote{``intl'' tokenizer, all data comes untokenized.} and ChrF~\citep{popovic-2015-chrf}.

\paragraph{Training and evaluation data} For our experiments, we trained diacritization models using data from three sources, which are, the \yoruba Bible, combination of training and development split of \menyo (now referred to as YAD training split from here onwards), and JW300~\citep{agic-vulic-2019-jw300}. \Cref{tab:data_count} shows the number of sentences in each split. Different combinations of these datasets were used for training our models, and unless specified otherwise we implemented the four previously described heuristics on the training data also. We evaluated our models on the YAD development and test splits, as well as on an existing Global Voices (GV) test set from~\cite{orife2020improving}. Additionally, we used a sampled $10\%$ of the Bible as a test set. For reproducibility, we release our code, data and models on GitHub\footnote{\url{https://github.com/ajesujoba/YAD}}.

\paragraph{Baseline  models and result} We trained diacritization models by fine-tuning MT5, Afri-MT5, AfriTeVa-V2, and UMT5, on the YAD training split. The experimental results as presented in \Cref{tab:main_result} shows that AfriTeVa-V2-large outperform other massively multilingual language models. Interestingly, Oyo-T5-base demonstrates competitiveness on both BLEU and CHRF similar to AfriTeVa-V2-large, despite having fewer parameters compared to AfriTeVa-V2-large. 

\begin{table}[htbp]
\scalebox{0.8}{%
\begin{minipage}[b]{0.4\linewidth}
    \centering
    \begin{tabular}{cccc}
        \toprule
        \textbf{Corpus} & \textbf{Train} & \textbf{Dev} & \textbf{Test} \\
        \midrule 
        Bible & 27739 & - & 3083 \\
        JW300 & 459,871 & - & - \\ 
        \rowcolor{GrayColor}
        YAD & 13,467 & 3,326 & 3,330 \\
        \midrule
        \multicolumn{3}{l}{Previous test set: \cite{orife2020improving}} \\
        \rowcolor{GrayColor}
        GV & - & - & 619 \\
        \bottomrule
    \end{tabular}
    \caption{Data split of different corpus.}
    \label{tab:data_count}
\end{minipage}}%
\scalebox{0.7}{%
\begin{minipage}[b]{1.0\linewidth}
    \centering
    \begin{tabular}{l|c|rrrr}
        \toprule
         &  & \multicolumn{2}{c}{\textbf{Dev. eval.}} & \multicolumn{2}{c}{\textbf{Test eval.}} \\
        \textbf{Model} & \textbf{Size} & \textbf{BLEU} & \textbf{CHRF} & \textbf{BLEU} & \textbf{CHRF} \\
        \midrule
        UMT5-base & 580M & 38.1 & 62.5 & 39.3 & 64.0 \\
        MT5-base & 580M & 48.1 & 69.4 & 49.5 & 71.2 \\
        AfriMT5-base & 580M & 53.3 & 72.8 & 54.0 & 74.2 \\
        AfriTeVa-base & 313M & 56.9 & 72.9 & 57.2 & 74.0 \\
        AfriTeVa-large & 871M & 67.2 & 80.8 & 66.8 & 81.3 \\
        Oyo-T5-base & 280M & \underline{\textbf{71.9}} & \underline{\textbf{82.2}} & \underline{\textbf{70.2}} & \underline{\textbf{82.3}} \\

        \bottomrule
    \end{tabular}
    \caption{Diacritization model evaluation on YAD Dev. and Test sets.}
    \label{tab:main_result}
\end{minipage}}
\end{table}

\paragraph{Effect of model scaling} Given that Oyo-T5-base is a competitive model despite its size, we chose to train light-weight diacritizers by finetuning the smaller variants of Oyo-T5 model on the YAD training data to check the effect of model size on this task. We evaluated the models on YAD dev. and test sets. As shown in \Cref{tab:model_scale}, bigger models are better for this task. However, Oyo-T5-small with 60M parameters even outperformed AfriTeVa-base with 313M parameters on the same task.

\paragraph{Effect of training data scaling} Going further, we choose to evaluate the effect of training data size on the diacritization models. Hence trained new diacritization models on Bible, JW300, and a combination of them and YAD training data using Oyo-T5-base as the backbone model and we evaluate them on the Bible, GV, and YAD test sets. For these test sets, we use the version where all diacritics were removed from the source side without applying the earlier described heuristics \footnote{We also analyze this effect on text with only underdot on the source side and provide the result in \Cref{sec:inputside}}.

\Cref{tab:data_scale} shows that training on more data is better. Also training on just JW300 outperformed the best result achieved by \citet{orife2020improving} on GV by $+11.6$ BLEU point. Furthermore, we observed that models trained on a specific domain tend to perform better on that domain. For instance, models trained on the Bible achieved a $+17.6$ BLEU point improvement compared to the next best performing model, which was trained on JW300. It is important to note that YAD's training data, obtained from the combination of \menyo training and development splits, included Global Voices news. Therefore, we suspect that its high BLEU scores on the GV tests set could be attributed to potential data leakage. Lastly, our findings suggest that more data generally leads to better performance, as evidenced by the results from the model trained on the combined dataset.

\begin{table}[htbp]
\begin{minipage}[t]{0.4\linewidth}
\centering

\scalebox{0.8}{%
\begin{tabular}{l|ccc}
\toprule
\textbf{Model/Data} & \textbf{Bible} & \textbf{GV} & \textbf{YAD} \\
\midrule
\multicolumn{4}{l}{Data: JW300+Bible+others} \\
\cite{orife2020improving}  & - & 59.8 & - \\
\midrule
Bible & 88.1 & 55.0 & 57.2 \\
JW300 & 71.7 & 71.4  & 66.9   \\
YAD & 69.0 & 82.1 & 72.0  \\
JW300+Bible+YAD & 90.8 & 91.0 & \textbf{78.7} \\ 
\bottomrule
\end{tabular}
}
\caption{Data scale effect on Bible, GV \& YAD.}
\label{tab:data_scale}
\end{minipage}%
\begin{minipage}[t]{0.45\linewidth}
\centering
\scalebox{0.89}{%
\begin{tabular}{l|c|cc|cc}
\toprule
 & & \multicolumn{2}{c|}{\textbf{Dev. eval.}} & \multicolumn{2}{c}{\textbf{Test eval.}} \\
\midrule
\textbf{Model} & \textbf{Size} & \textbf{BLEU} & \textbf{CHRF} & \textbf{BLEU} & \textbf{CHRF} \\
\midrule
Oyo-T5-tiny & 14M & 42.1 & 62.1 & 36.3 & 60.4 \\
Oyo-T5-mini & 18M & 49.0 & 64.6 & 47.5 & 65.6 \\
Oyo-T5-small & 60M & 63.5 & 75.8 & 62.9 & 76.7 \\
Oyo-T5-base & 280M & \underline{\textbf{71.9}} & \underline{\textbf{82.2}} & \underline{\textbf{70.2}} & \underline{\textbf{82.3}} \\ 
\bottomrule
\end{tabular}
}
\caption{Effect of the model scale on YAD.\looseness-1}
\label{tab:model_scale}
\end{minipage}
\end{table}

\paragraph{Conclusion} In this work we present \yoruba automatic diacritization (YAD) benchmark dataset for evaluating \yoruba diacritization systems. In addition, we pre-train text-to-text transformer, T5 model for \yoruba and showed that this model outperform several multilingually trained T5 models. Lastly, we showed that more data and bigger models are better at diacritization for \yoruba.

\subsubsection*{Acknowledgments}
Jesujoba Alabi was partially supported by the BMBF’s (German Federal Ministry of Education and Research) SLIK project under the grant 01IS22015C. David Adelani acknowledges the support of DeepMind Academic Fellowship programme. This work was supported in part by Oracle Cloud credits and related resources provided by Oracle.

\bibliography{iclr2024_conference}
\bibliographystyle{iclr2024_conference}

\appendix
\section{Effect of applying diacritics to undiacritized texts versus text with only underdot}
\label{sec:inputside}
\begin{table}[h]
\centering
\scalebox{0.9}{%
\begin{tabular}{l|cc|cc|cc}
\toprule
 & \multicolumn{2}{c|}{\textbf{Bible}} & \multicolumn{2}{c|}{\textbf{GV.}} & \multicolumn{2}{c}{\textbf{YAD.}}\\
\midrule
\textbf{Model} & \textbf{BLEU} & \textbf{CHRF} & \textbf{BLEU} & \textbf{CHRF}& \textbf{BLEU} & \textbf{CHRF} \\
\midrule
\multicolumn{7}{c}{\textbf{no accent on source side}} \\
Bible &  88.2 & 94.3 & 55.2 & 68.4 & 57.8 & 72.9 \\
JW300 &  64.0 & 80.9 & 68.8 & 81.6 & 65.0 & 80.9 \\
Menyo &  71.8 & 83.9 & 79.3 & 86.7 & 72.0 & 83.7 \\
JW300+Bible+Menyo & 89.6 & 95.8 & 90.0 & 95.1 & 78.4 & 89.0 \\ 
\midrule
\multicolumn{7}{c}{\textbf{no diacritics on source side}} \\
Bible &  88.1 & 94.2 & 55.0 & 67.4 & 57.2 & 71.6 \\
JW300 &  71.7 & 82.6 & 71.4 & 83.7 & 66.9 & 81.9 \\
Menyo &  69.0 & 82.4 & 82.1 & 87.3 & 72.0 & 83.3 \\
JW300+Bible+Menyo & 90.8 & 96.0 & 91.0 & 95.1 & 78.7 & 88.8 \\ 
\bottomrule
\end{tabular}
}
\caption{Effect of applying diacritics to undiacritized texts vs text with only underdots}
\label{tab:model_scale2}
\end{table}

\end{document}